# Learning Patient-Specific Disease Dynamics with Latent Flow Matching for Longitudinal Imaging Generation

### v2

@@ 4.2 Benchmark Comparisons
We present qualitative results in Figure 2. In this subject, disease progression primarily affects the lateral ventricles (the central butterfly-shaped region) and the cortical gray matter (outer surface). SADM, TADM, and MambaControl fail to capture changes in the lateral ventricles, resulting in pronounced butterfly-shaped error in the residuals image. DiffuseMorph and BrLP better capture ventricular changes, but leave substantial residuals along the cortical surface. Our method also shows error in the ventricles; however, these are less intense, with thinner boundaries and smaller cortical-surface errors, demonstrating improved overall accuracy.
@@ 4.3 Ablation study
We provide visualization results of the learned latent space in Figure 3 over test cases in ADNI dataset, where t-SNE is used to project the representations into two dimensions. As shown in the left panel, the latent trajectories of individual patients are well aligned. Interestingly, the right panel reveals that the latent space also organizes according to diagnosis status, despite the fact that this information was never used during training.
@@
The baseline Latent Flow Matching (LFM) (dao2023flow) model without conditioning performs the worst across all metrics, showing notably low SSIM/PSNR and high error values. Incorporating conditional information (Conditional LFM) significantly improves image quality (SSIM from 88.62 to 91.20; PSNR from 27.59 to 28.46) and reduces both Region MAE and $\Delta$-RMAE.
@@ 5 Conclusion
We introduce Progression Latent Flow Matching ($\Delta$-LFM), a framework that unifies the proposed ArcRank loss with temporal flow matching to capture patient-specific disease trajectories. To evaluate progression in isolation, we propose a new metric, $\Delta$-RMAE. On longitudinal MRI benchmarks, $\Delta$-LFM consistently achieves higher fidelity, lower error, and tighter alignment with ground-truth progression than prior methods. Beyond accuracy, it yields temporally coherent and clinically interpretable trajectories. These results show that the $\Delta$-LFM provides a new direction for progression modeling, with the potential to generalize across other temporal generation tasks.
@@ Appendix A Training Setting
We adopt the AdamW optimizer (loshchilov2017decoupled) for all experiments. For the autoencoder (AE) and 3D U-Net (ronneberger2015u), we trained with a learning rate of $1\times 10^{-3}$ and a batch size of $2$ for 300 epochs. For the flow matching model based on the 3D U-Net, we used a reduced learning rate of $3\times 10^{-5}$ with a larger batch size of $4$ for 200 epochs.
@@ Appendix B Conditional Generation
Following Litrico et al. (litrico2024tadm), we encode the current sample time in flow matching into continuous embeddings using sinusoidal positional encodings. These time embeddings act as an additive bias to the input, ensuring that the generative process is aware of its temporal position.
Building on this, and inspired by BrLP (puglisi2024enhancing), we introduce explicit control signals that combine patient-specific attributes with temporal queries. In particular, we condition the model not only on the start time, the current sample time, and the end query time, but also on the patient attributes. This design enables the model to incorporate both subject-specific and task-specific information when shaping the generative trajectory.
To inject these conditional signals into the U-Net backbone, we explored three strategies: 1. Additive biasing, 2. Cross-attention (vaswani2017attention) and 3. Adaptive Layer Normalization (AdaLN) (raffel2020exploring).
@@ Appendix C Auto-encoder Ablation
We do not use deeper or wider AEs, as these configurations lead to out-of-memory (OOM) errors on our RTX A6000 GPU (48 GB). Based on the above results, we use channels $[64,128,256]$ and crop size $64^{3}$ as our default configuration. Larger crops (e.g., $72^{3}$) also result in OOM on our hardware. To keep the comparison with baselines fair, we do not use a pretrained AE, although initializing with MAISI guo2025maisi can further improve PSNR (from 30.04 to 31.97 for $[64,128,256]$).
@@ Appendix D Longitudinal Progression Quantitative Results
Appendix E On the Stability of $\Delta$-RMAE
@@ Appendix I Discussion of LLM usage
We used large language models (LLMs), including ChatGPT (openai2023chatgpt) and Gemini (team2023gemini), as writing assistants for grammar correction and style refinement. All technical content, methodology, and experiments were developed by the authors.

### v3

@@ 4.2 Benchmark Comparisons
We present qualitative results in Figure 2. In this subject, disease progression primarily affects the lateral ventricles (the central butterfly-shaped region) and the cortical gray matter (outer surface). SADM, TADM, and MambaControl fail to capture changes in the lateral ventricles, resulting in pronounced butterfly-shaped error in the residual image. DiffuseMorph and BrLP better capture ventricular changes, but leave substantial residuals along the cortical surface. Our method also shows error in the ventricles; however, these are less intense, with thinner boundaries and smaller cortical-surface errors, demonstrating improved overall accuracy.
@@ 4.3 Ablation study
We provide visualization results of the learned latent space in Figure 3 on test cases in ADNI dataset, where t-SNE is used to project the representations into two dimensions. As shown in the left panel, the latent trajectories of individual patients are well aligned. Interestingly, the right panel reveals that the latent space also organizes according to diagnostic status, despite the fact that this information was never used during training.
@@
The baseline Latent Flow Matching (LFM) (Dao et al., 2023) model without conditioning performs the worst across all metrics, showing notably low SSIM/PSNR and high error values. Incorporating conditional information (Conditional LFM) significantly improves image quality (SSIM from 88.62 to 91.20; PSNR from 27.59 to 28.46) and reduces both Region MAE and $\Delta$-RMAE.
@@ 5 Conclusion
We introduce Progression Latent Flow Matching ($\Delta$-LFM), a framework that unifies the proposed ArcRank loss with temporal flow matching to capture patient-specific disease trajectories. To evaluate progression in isolation, we propose a new metric, $\Delta$-RMAE. On longitudinal MRI benchmarks, $\Delta$-LFM consistently achieves higher fidelity, lower error, and tighter alignment with ground-truth progression than prior methods. Beyond accuracy, it yields temporally coherent and clinically interpretable trajectories. These results show that $\Delta$-LFM provides a new direction for progression modeling, with the potential to generalize across other temporal generation tasks.
@@ Appendix A Training Setting
We adopt the AdamW optimizer (Loshchilov & Hutter, 2017) for all experiments. For the autoencoder (AE) and 3D U-Net (Ronneberger et al., 2015), we trained with a learning rate of $1\times 10^{-3}$ and a batch size of $2$ for 300 epochs. For the flow matching model based on the 3D U-Net, we used a reduced learning rate of $3\times 10^{-5}$ with a larger batch size of $4$ for 200 epochs.
@@ Appendix B Conditional Generation
Following Litrico et al. (Litrico et al., 2024), we encode the current sample time in flow matching into continuous embeddings using sinusoidal positional encodings. These time embeddings act as an additive bias to the input, ensuring that the generative process is aware of its temporal position.
Building on this, and inspired by BrLP (Puglisi et al., 2024), we introduce explicit control signals that combine patient-specific attributes with temporal queries. In particular, we condition the model not only on the start time, the current sample time, and the end query time, but also on the patient attributes. This design enables the model to incorporate both subject-specific and task-specific information when shaping the generative trajectory.
To inject these conditional signals into the U-Net backbone, we explored three strategies: (1) Additive biasing, (2) Cross-attention (Vaswani et al., 2017) and (3) Adaptive Layer Normalization (AdaLN) (Raffel et al., 2020).
@@ Appendix C Auto-encoder Ablation
We do not use deeper or wider AEs, as these configurations lead to out-of-memory (OOM) errors on our RTX A6000 GPU (48 GB). Based on the above results, we use channels $[64,128,256]$ and crop size $64^{3}$ as our default configuration. Larger crops (e.g., $72^{3}$) also result in OOM on our hardware. To keep the comparison with baselines fair, we do not use a pretrained AE, although initializing with MAISI (Guo et al., 2025) can further improve PSNR (from 30.04 to 31.97 for $[64,128,256]$).
@@ Appendix D Longitudinal Progression Quantitative Results
Appendix E Stability of $\Delta$-RMAE
@@ Appendix I Discussion of LLM usage
We used large language models (LLMs), including ChatGPT (OpenAI, 2023) and Gemini (Team et al., 2023), as writing assistants for grammar correction and style refinement. All technical content, methodology, and experiments were developed by the authors.